# AI-Generated Images Introduce Invisible Relevance Bias to Text-Image Retrieval

### 23 Nov 2023 v1

@@ 2.2.2. Human Evaluation
To further assess whether the AI-generated images in our constructed benchmark introduce additional visual semantics that are more relevant to the queries in retrieval, we conduct the human evaluation. Specifically, we invite five volunteers with master’s degrees to participate in the human evaluation. The triple consisting of the caption (i.e., the query in retrieval), real image, and its corresponding AI-generated image is the basic unit in this evaluation. We ask the five volunteers to select which image (real or AI-generated) has more relevant semantics to the caption. If the volunteers think the two images are equally relevant or irrelevant to the caption, they can select "equal". We count the proportion of selections made by volunteers on the entire benchmark and the results are shown in Table 3. Volunteers in human evaluation think that in most samples of our benchmark, real images have more or equal visual semantics relevant to the caption (i.e., the query in retrieval) than AI-generated images. This result further guarantees the fairness of our assessment of invisible relevance bias. If the AI-generated images do not have more relevant visual semantics than real images, while the text-image retrieval model still tends to rank AI-generated images higher than the real images, the invisible relevance bias does exist.
Table 3. Proportion of the selections made by volunteers in human evaluation.
Flicker30k+AIGC MSCOCO+AIGC
@@ 3. Bias Assessment
Based on the benchmark we constructed above, we can assess the impact of AI-generated images on text-image retrieval models. We not only study the impact on ranking results when AI-generated images appear in the retrieved database but also further discuss the problems caused when retrieval models are trained on the training data mixed with AI-generated images.
@@ 3.1. Text-Image Retrieval Models
From the perspective of the training data of models, our assessment includes not only the retrieval models trained from scratch on supervised text-image pairs but also vision-language models that have been pre-trained on massive image-text pairs. From the perspective of the architecture of models, our assessment includes both dual-encoder-based and fusion-based models. We do not select models that are not open source because of the relatively high implementation costs. The specific models used in the assessment include: (1) NAAF (Zhang et al., 2022) is a fusion encoder text-image retrieval model that exploits both the positive effect of matched fragments and the negative effect of mismatched fragments to jointly infer text-image similarity. (2) VSE (Chen et al., 2021) is a dual-encoder text-image retrieval model that learns to automatically adapt the best pooling strategy for visual semantic embedding. We following the papers to train NAAF and VSE on supervised text-image retrieval datasets. (3) VILT is a fusion encoder text-image matching model based on the interaction between image and text in Transformer (Vaswani et al., 2017). It is a vision-language model that has been pre-trained on massive text-image pairs. (4) FLAVA is a language and vision alignment model that learns strong representations from multimodal. It is a vision-language model pre-trained on massive text-image pairs and can be used as the dual encoder model for text-image retrieval. (5) ALIGIN is also a dual encoder vision-language that has been pre-trained on over one billion image-text pairs. (6) BEIT-3 is a general-purpose multimodal foundation model that has been pre-trained on hundreds of millions of text-image pairs and massive texts and images. It can be used as a dual-encoder text-image retrieval model.
3.2. Experimental Settings and Metrics
As the neural networks adapt not only to the features of the task signal but also to the data domain during the training process (Xu et al., 2022), our assessment is performed under the out-of-domain setting to try to mitigate potential bias introduced by the domain of the training data. Specifically, for the models that need to train from scratch on supervised text-image pairs (NAAF and VSE), we train them on dataset A and evaluate their performance on another dataset B. For the models that have been pre-trained on massive real text-image pairs and show excellent zero-shot performance in text-image retrieval, we directly use these pre-trained models to perform retrieval on the test datasets. An exception is that even though VILT has been pre-trained, it needs to be combined with a specific multi-layer perceptron to complete the text-image matching task in text-image retrieval. So we fine-tune VILT on supervised text-image retrieval datasets just like NAAF and VSE.
The main metrics in this section aim to measure the difference between the ranking of real and AI-generated images in the retrieved results. We follow (Dai et al., 2023) to measure the relative percentage difference between them as:
in which Metric can be the evaluation metrics for information retrieval such as NDCG@k and R@k. $\rm{Relevance\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relevance\triangle}<0$ means retrieval models rank AI-generated images higher than real images. The absolute value of $\rm{Relevance\triangle}$ indicates the value of this bias (Dai et al., 2023).
3.3. Invisible Relevance Bias
Table 4. Performance of the retrieval models on the benchmark we constructed consisting of both real and AI-generated images. $\rm{Relevance\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relevance\triangle}<0$ means retrieval models rank AI-generated images higher than real images. The absolute value of $\rm{Relevance\triangle}$ indicates the degree of the bias.
Flicker30k+AIGC MSCOCO+AIGC
@@ 3.3. Invisible Relevance Bias
The performance of text-image retrieval models on the benchmark we constructed consisting of both real and AI-generated images are shown in Table 4. Overall, invisible relevance bias widely exists in text-image retrieval models, that is, text-image retrieval models tend to rank AI-generated images higher than real images even though they have very similar semantics. Specifically, the following conclusions can be made: (1) The invisible relevance bias exists in both the models trained from scratch and the vision-language models that have been pre-trained on massive supervised text-image pairs. (2) The invisible relevance bias exists in both dual-encoder-based and fusion-encoder-based retrieval models. (3) The invisible relevance bias has a relatively greater impact on the Top-1 retrieved image. In the retrieved list, the Top-1 item is most likely to be clicked by users, which means that invisible relevance bias introduced by AI-generated images will have a huge impact on users’ actual search and click results.
@@ 3.4. More Serious Bias Caused by Training
(b) $\rm{Relative}\triangle$ on NDCG@k of Flicker30k
(c) NDCG@k on only real or AI-generated images of Flicker30k
@@
(f) NDCG@k on only real or AI-generated images of MSCOCO
Figure 3. In-domain and out-of-domain assessment results of VSE trained on the datasets that reconstructed based on the training set of Flicker30k. We change the ratio of AI-generated images in the datasets (X-axis) while keeping the total number of samples in the training set unchanged. The model is tested on the test set of Flicker30k+AIGC (in-domain) and MSCOCO+AIGC (out-of-domain) respectively that we constructed in Section 2.
@@
Figure 4. Distribution of the relevance scores between captions and images estimated by retrieval models trained on the datasets mixed with different ratios of AI-generated images. Flicker30k is an in-domain setting and MSCOCO is an out-of-domain setting.
An essential point that needs to be noticed is that AI-generated images do not only have an impact on the inference stage of the retrieval models. Due to their large-scale distribution on the Internet and the bias that has been found in Section 3.3, AI-generated images are very likely to be mixed into the training data of retrieval models in the future. In this section, we delve deeper into the repercussions on retrieval performance and invisible relevance bias when AI-generated images are mixed into the training of retrieval models. Specifically, we explore the impact on the retrieval models when the training set is mixed with generated images by changing the ratio between real and generated images in the supervised text-image training set. To ensure a comprehensive and accurate assessment, our experiments focus on the model trained from scratch (VSE). This is because pre-trained vision-language models have been trained on a large number of real text-image pairs, which may introduce biases in evaluating the impact of different ratios of generated images in the training data. In this experiment, we reconstruct the training set of Flicker30k by replacing a certain proportion of real images with AI-generated images. Specifically, for each real image ($I_{r}$) in the training set of Flicker30k, we use the method in Section 2 to generate its corresponding AI-generated image ($I_{g}$). Then, we introduce the ratio $\alpha$, which means that in our reconstructed training data, the paired images for $\alpha$ percentage of captions are AI-generated images, and for $(100-\alpha)$ percentage of captions are real images. We change the ratio while keeping the total number of samples in the training set unchanged. We evaluate the performance of the trained model in the test set of Flicker30k+AIGC (in-domain setting) and MSCOCO+AIGC (out-of-domain setting).
The experimental results are shown in Figure 3 and Figure 4. Figure 3 illustrates that as the ratio increases, the ranking disparity between generated images and real images widens, with the retrieval model exhibiting a greater inclination to rank generated images higher. Meanwhile, Figure 4 demonstrates that with an increasing ratio, the mean of the score distribution for generated images rises. Simultaneously, the difference in score distribution between generated images and real images grows, and the scores of AI-generated images gradually become greater and greater than those of real images. The following conclusions can be drawn from these results: (1) Retrieval models trained on the datasets mixed with AI-generated images exhibit more serious invisible relevance bias in both in-domain and out-of-domain settings. (2) The invisible relevance bias tends to become more obvious as the ratio of AI-generated images in the training datasets increases in both in-domain and out-of-domain settings. (3) As the ratio of AI-generated images in the training datasets increases, the retrieval performance on real images is gradually damaged. This phenomenon is more obvious in the in-domain setting.
The experimental results in this section also reveal a vicious cycle of falling into invisible relevance bias. Specifically, the invisible relevance bias of text-image retrieval models causes the AI-generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of retrieval models, leading to more serious bias and forming a vicious cycle. This ultimately results in users’ search results being surrounded by AI-generated images. This indicates that attention should be paid to alleviating the ranking bias caused by AI-generated images.
@@ 4. Our Method in debiasing
Section 3 reveals that invisible relevance bias widely exists in different text-image retrieval models and the training on data mixed with AI-generated images makes this bias more serious, which forms a vicious cycle and results in users’ search results being surrounded by AI-generated images. Therefore, it is very important to propose a method to mitigate this bias. In this section, we propose an effective method to achieve it by measuring and optimizing the invisible relevance bias in training. The specific details of our method will be introduced in the following contents.
@@ 4.1. Design of Debiasing Method
Our method alleviates the retrieval model’s preference bias for generated images by measuring and optimizing the additionally introduced invisible relevance in the training. For a caption-image pair $(C_{i},I^{r}_{i})$ in the training set, whether the dual-encoder or fusion-encoder architecture text-image retrieval models, the training process can be summarized as estimating the relevance score between $C_{i}$ and $I^{r}_{i}$, and using contrastive loss or regression loss as the optimization objective, which can be described as:
$s$ $=R(C_{i},I^{r}_{i};\theta),$
$\theta_{\text{optimal}}$ $=\arg\min_{\theta}\mathcal{L}(s,y;\theta),$
in which $R(\cdot;\theta)$ is the relevance scoring function of the retrieval model such as cosine similarity between embeddings in the dual-encoder model and neural networks in the fusion-encoder model, $\theta$ is the set of parameters of the model, $\mathcal{L}$ is the loss function such as contrastive loss or regression loss, $y$ is the label. In our method, for each real image $I^{r}_{i}$ in the training data, we use the method in Section 2 to generate its corresponding AI-generated image $I^{g}_{i}$. Then we can get the caption-real-AI triple as $(C_{i},I^{r}_{i},I^{g}_{i})$, in which $C_{i}$ is the paired caption for the real image $I^{r}_{i}$. We introduce the contrastive loss to get the difference in relevance scores between $I^{r}_{i}$ and $I^{g}_{i}$ with respect to the caption $C_{i}$ as:
@@
4.2. Evaluation of Debiasing
Table 5. Performance of the retrieval models on the benchmark we constructed consisting of both real and AI-generated images with different sampling probability $\beta$ in our debiasing method. $\rm{Relevance\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relevance\triangle}<0$ means retrieval models rank AI-generated images higher than real images. When $\rm{Relevance\triangle}<0$, the absolute value of $\rm{Relevance\triangle}$ indicates the value of this bias.
Flicker30k+AIGC (In-domain) MSCOCO+AIGC (Out-of-domain)
$\beta$ $\beta$
w/o debias 50% 60% 70% 80% 90% 100% w/o debias 50% 60% 70% 80% 90% 100%
@@ 4.2. Evaluation of Debiasing
Figure 5. Distribution of the relevance scores between captions and images estimated by retrieval models with different sampling probability $\beta$ in our debiasing method. Flicker30k is an in-domain setting and MSCOCO is an out-of-domain setting.
(a) $\beta=50\%$ on Flicker30k
(b) $\beta=60\%$ on Flicker30k
(c) $\beta=70\%$ on Flicker30k
(d) $\beta=80\%$ on Flicker30k
(e) $\beta=90\%$ on Flicker30k
(f) $\beta=100\%$ on Flicker30k
(g) $\beta=50\%$ on MSCOCO
(h) $\beta=60\%$ on MSCOCO
(i) $\beta=70\%$ on MSCOCO
(j) $\beta=80\%$ on MSCOCO
(k) $\beta=90\%$ on MSCOCO
(l) $\beta=100\%$ on MSCOCO
@@
We evaluate the effect of our debiasing method from three perspectives: (1) How our method affects the ranking difference between real images and AI-generated images. (2) How our method affects the distribution of relevance scores between captions and images. (3) How our method affects the representations of real and AI-generated images encoded by the image encoder of the retrieval model. All of these three evaluations are performed in both in-domain and out-of-domain settings, which can provide a comprehensive evaluation of our proposed method.
Table 5 shows how the $\rm{Relative\triangle}$ and retrieval performance changes with the sampling probability $\beta$. The results indicate that: (1) Our method not only effectively alleviates the retrieval model’s preference for AI-generated images, but also makes real images ranked significantly higher than AI-generated images in the final ranked list. (2) As the sampling probability $\beta$ increases, the retrieval model can better avoid the interference of AI-generated images and better expose real images to users. (3) When $\beta$ is $0.5$, retrieval models can achieve a fair ranking between real and generated images with little bias. (4) Our method not only does not impair the retrieval performance but also enhances it. It is mainly because, under the constraints of our method, AI-generated images can be seen as the hard negatives in training, which enhances the ability of the retrieval model to estimate the relevance, especially for distinguishing the images with very similar semantics.
Figure 5 shows the distribution of the relevance scores between captions and images estimated by retrieval models with different sampling probability $\beta$. The results indicate that our method effectively reduces the relevance scores to AI-generated images estimated by the retrieval model. With an increase in the sampling probability ($\beta$), the disparity in score distribution between the generated images and the real images expands, and the scores of real images gradually become greater and greater than those of AI-generated images.
Figure 6 shows the T-SNE visualization of the embeddings of real and AI-generated images acquired from the image encoder in VSE. The findings reveal that, with an increase in the sampling probability ($\beta$), the representations of generated and real images become more distinguishable. This observation indicates that our method empowers the retrieval model to capture the feature differences between generated and real images, reflecting these distinctions in the representations.

### 15 Jan 2024 v3

@@ 2.2.2. Human Evaluation
Table 3 shows humans think that in most samples of our benchmark, real images have more or equal visual semantics relevant to the queries than AI-generated images. This further guarantees the fairness of our assessment of invisible relevance bias. If the AI-generated images do not have more relevant visual semantics than real images, while the text-image retrieval model still tends to rank AI-generated images higher than the real images, the invisible relevance bias does exist. Specifically, we invite five humans with master’s degrees to participate in the evaluation. The triple consisting of a caption (i.e., the query), a real image, and its corresponding AI-generated image is the basic unit in this evaluation. We ask the five humans to select which image (real or AI-generated) has more relevant semantics to the caption. We count the proportion of selections made by humans on our benchmark and the results are shown in Table 3.
Table 3. Proportion of the selections made by humans.
Flicker30k+AI MSCOCO+AI
@@ 3. Bias Assessment
Based on the benchmark constructed above, we can assess the impact of AI-generated images on text-image retrieval. We not only study the impact when AI-generated images appear in the database but also further discuss the problems caused when AI-generated images are mixed into the training data of retrieval models.
@@ 3.1. Text-Image Retrieval Models
As for training data, our assessment includes not only the retrieval models trained from scratch on supervised text-image pairs but also vision-language models that have been pre-trained on massive image-text pairs. As for model architecture, our assessment includes both dual-encoder-based and fusion-based models. The specific models used in the assessment include: (1) NAAF (Zhang et al., 2022) is a fusion encoder text-image retrieval model that exploits both the positive effect of matched fragments and the negative effect of mismatched fragments to jointly infer text-image similarity. (2) VSE (Chen et al., 2021) is a dual-encoder text-image retrieval model that learns to automatically adapt the best pooling strategy for visual semantic embedding. (3) VILT is a fusion encoder text-image matching model based on the interaction between image and text in Transformer (Vaswani et al., 2017). It is a vision-language model that has been pre-trained on massive text-image pairs. (4) FLAVA learns strong representations from multimodal. It is a vision-language model pre-trained on massive text-image pairs and can be used as the dual encoder model for text-image retrieval. (5) ALIGIN is also a dual encoder vision-language that has been pre-trained on over one billion image-text pairs. (6) BEIT-3 is a multimodal foundation model that has been pre-trained on hundreds of millions of text-image pairs and massive texts and images. It can be used as a dual-encoder text-image retrieval model.
Table 4. Performance of the retrieval models on the benchmark we constructed consisting of both real and AI-generated images. $\rm{Relative\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relative\triangle}<0$ means retrieval models rank AI-generated images higher than real images. The absolute value of $\rm{Relative\triangle}$ indicates the degree of the bias.
Flicker30k+AI MSCOCO+AI
@@
3.2. Experimental Settings and Metrics
As the neural networks tend to fit the data domian in training (Xu et al., 2022), our assessment is performed under the out-of-domain setting to try to mitigate potential bias introduced by the domain of the training data. Specifically, for the models that need to train from scratch on supervised text-image pairs (NAAF and VSE), we train them on Flicker30k (MSCOCO) and evaluate their performance on MSCOCO (Flicker30k). For the models that have been pre-trained on massive real text-image pairs and show excellent zero-shot performance in text-image retrieval, we directly use these pre-trained models to perform retrieval on the test datasets. An exception is that even though VILT has been pre-trained, it needs to be combined with a specific multi-layer perceptron to complete the text-image matching task in text-image retrieval. So we fine-tune VILT on supervised text-image retrieval datasets just like NAAF and VSE.
The metric follows (Dai et al., 2023) to measure the difference between the ranking of real and AI-generated images in the retrieved results as:
in which Metric can be the metrics for IR such as NDCG@k and R@k. $\rm{Relative\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relative\triangle}<0$ means retrieval models rank AI-generated images higher than real images. The absolute value of $\rm{Relative\triangle}$ indicates the degree of the bias (Dai et al., 2023).
3.3. Invisible Relevance Bias
The experimental results are shown in Table 4. Overall, invisible relevance bias widely exists in text-image retrieval models, that is, text-image retrieval models tend to rank AI-generated images higher than real images even though they have very similar semantics. Specifically, the following conclusions can be made: (1) The invisible relevance bias exists in both the models trained from scratch and the vision-language models that have been pre-trained on massive supervised text-image pairs. (2) The invisible relevance bias exists in both dual-encoder-based and fusion-encoder-based retrieval models. (3) The invisible relevance bias has a relatively greater impact on the Top-1 retrieved image. In the retrieved list, the Top-1 item is most likely to be clicked by users, which means that invisible relevance bias introduced by AI-generated images will have a huge impact on users’ actual search and click results.
@@ 3.4. More Serious Bias Caused by Training
Due to wide distribution of AI-generated images on the Internet and the bias in Section 3.3, AI-generated images are very likely to be mixed into the training data of retrieval models. This section delves deeper into the impact on retrieval performance and invisible relevance bias when AI-generated images are mixed into the training of retrieval models. The experimental results reveal a vicious cycle of falling into invisible relevance bias. Specifically, the invisible relevance bias of text-image retrieval models causes the AI-generated images to have a higher chance of being obtained from massive data, which makes them more likely to be mixed into the training of retrieval models, leading to more serious bias and forming a vicious cycle. This ultimately results in users’ search results being surrounded by AI-generated images.
Specifically, we explore the impact of the training mixed with generated images on retrieval by incorporating varying ratios of generated images into the training set. To ensure an accurate assessment, our experiments focus on the model trained from scratch (VSE). This is because pre-trained vision-language models have been pre-trained on massive real text-image pairs, which introduce biases in evaluating the impact of different ratios of generated images in the training data. We reconstruct the training set of Flicker30k by replacing a certain ratio of real images with AI-generated images. For each real image ($I_{r}$) in the training set, we use the method in Section 2 to generate its corresponding AI-generated image ($I_{g}$). Then, we introduce the ratio $\alpha$, which means that in our reconstructed training data, the paired images for $\alpha$ percentage of captions are AI-generated images, and for $(100-\alpha)$ percentage of captions are real images. We change the ratio while keeping the total number of training samples unchanged. We evaluate the performance of the trained model on the test set of Flicker30k+AI (in-domain) and MSCOCO+AI (out-of-domain).
@@
(b) $\rm{Relative}\triangle$ on NDCG of Flicker30k
(c) NDCG on only real or AIGC images of Flicker30k
@@
(f) NDCG on only real or AIGC images of MSCOCO
Figure 2. Assessment results on the training set miexed with AI-generated images. We change the ratio of AI-generated images in the datasets (X-axis) while keeping the total number of training samples unchanged. The model is tested on the test set of Flicker30k+AI (in-domain) and MSCOCO+AI (out-of-domain) respectively that we constructed in Section 2.
@@
Figure 3. Distribution of the caption-image relevance scores estimated by retrieval models that are trained on the datasets mixed with different ratios of AI-generated images. Flicker30k is in-domain and MSCOCO is out-of-domain.
Figure 2 illustrates that as the ratio increases, the ranking disparity between generated images and real images widens, with the retrieval model exhibiting a greater inclination to rank generated images higher. Meanwhile, Figure 3 demonstrates that with an increasing ratio, the discrepancy in score distribution between generated images and real images increases, and the scores of AI-generated images gradually become greater and greater than those of real images. In both in-domain and out-of-domain settings, the following conclusions can be drawn from these results: (1) Retrieval models trained on the datasets mixed with AI-generated images exhibit more serious invisible relevance bias. (2) The invisible relevance bias tends to become more obvious as the ratio of AI-generated images in the training datasets increases. (3) As the ratio of AI-generated images in the training datasets increases, the retrieval performance on real images is gradually damaged.
@@ 4. Our Method in debiasing
In this section, we propose an effective method to alleviate the invisible relevance bias. The details will be introduced in the following.
@@ 4.1. Design of Debiasing Method
Our method alleviates the retrieval model’s preference bias for generated images by measuring and optimizing the additionally introduced invisible relevance in the training. Given a caption-image pair $(C_{i},I^{r}_{i})$ in the training set, the training process for both dual-encoder and fusion-encoder text-image retrieval models can be formulated as estimating the relevance score $s$ between $C_{i}$ and $I^{r}_{i}$, and using contrastive loss or regression loss as the optimization objective to adjust $s$, which can be described as:
$s=R(C_{i},I^{r}_{i};\theta),\quad\theta_{\text{optimal}}=\arg\min_{\theta}\mathcal{L}(s,y;\theta),$
in which $R(\cdot;\theta)$ is the relevance scoring function of the retrieval model such as cosine similarity between representations in the dual-encoder model and neural networks in the fusion-encoder model, $\theta$ is the set of parameters of the model, $\mathcal{L}$ is the loss function such as contrastive loss or regression loss, $y$ is the label. In our method, for each real image $I^{r}_{i}$ in the training data, we use the method in Section 2 to generate its corresponding AI-generated image $I^{g}_{i}$. Then we can get the caption-real-AI triple as $(C_{i},I^{r}_{i},I^{g}_{i})$, in which $C_{i}$ is the paired caption for the real image $I^{r}_{i}$. We introduce the contrastive loss to get the difference in relevance scores between $I^{r}_{i}$ and $I^{g}_{i}$ with respect to the caption $C_{i}$ as:
@@
Table 5. Performance of the retrieval models on the benchmark we constructed consisting of both real and AI-generated images with different sampling probability $\beta$ in our debiasing method. $\rm{Relative\triangle}>0$ means retrieval models rank real images higher than AI-generated images, $\rm{Relative\triangle}<0$ means retrieval models rank AI-generated images higher than real images. When $\rm{Relative\triangle}<0$, the absolute value of $\rm{Relative\triangle}$ indicates the value of this bias.
Flicker30k+AI (In-domain) MSCOCO+AI (Out-of-domain)
w/o debias $\beta=50\%$ $\beta=60\%$ $\beta=70\%$ $\beta=80\%$ $\beta=90\%$ $\beta=100\%$ w/o debias $\beta=50\%$ $\beta=60\%$ $\beta=70\%$ $\beta=80\%$ $\beta=90\%$ $\beta=100\%$
@@
Figure 4. Distribution of the caption-image relevance scores estimated by retrieval models with different sampling probability $\beta$ in our debiasing method. Flicker30k is in-domain and MSCOCO is out-of-domain.
(a) $\beta=60\%$ on Flicker30k
(b) $\beta=80\%$ on Flicker30k
(c) $\beta=100\%$ on Flicker30k
(d) $\beta=60\%$ on MSCOCO
(e) $\beta=80\%$ on MSCOCO
(f) $\beta=100\%$ on MSCOCO
@@
4.2. Evaluation of Debiasing
We evaluate the effect of our debiasing method from three perspectives: (1) How our method affects the ranking difference between real images and AI-generated images. (2) How our method affects the distribution of caption-image relevance scores. (3) How our method affects the representations of real and AI-generated images. All of these three evaluations are performed in both in-domain (Flicker30k) and out-of-domain (MSCOCO) setting.
Table 5 shows how the $\rm{Relative\triangle}$ and retrieval performance changes with the sampling probability $\beta$. The results indicate that: (1) Our method not only effectively alleviates the retrieval model’s preference for AI-generated images, but also makes real images ranked significantly higher than AI-generated images. (2) As the sampling probability $\beta$ increases, real images are ranked higher and higher than AI-generated images. (3) When $\beta$ is $0.5$, retrieval models can achieve a fair ranking between real and generated images with little bias. (4) Our method improves the retrieval performance. It is because in our method, AI-generated images can be seen as the hard negatives in training, which enhances the ability of the retrieval model to distinguish the images with very similar semantics.
Figure 4 shows the distribution of the caption-image relevance scores estimated by retrieval models with different sampling probability $\beta$. The results indicate that our method effectively reduces the relevance scores between captions and generated images estimated by the retrieval model. With the increase in sampling probability ($\beta$), the disparity in score distribution between the generated images and the real images expands, and the scores of real images gradually become greater and greater than those of generated images.
Figure 5 shows the T-SNE of the embeddings of real and AI-generated images encoded by image encoder in VSE. The reveals that with the increase in the sampling probability ($\beta$), the representations of generated and real images become more distinguishable.